ln x + log x = 4 (solve for x to the 3rd decimal)
Answer by Alan3354(69443)

ln x + log x = 4
log(x)*ln(10) + log(x) = 4
log(x)*(ln(10 + 1)) = 4
log(x) = 4/(ln(10)+1)
log(x) =~ 1.211172
x =~ 16.262